OpenStudy (anonymous):

factor completely a^2-2ab-15b^2 how do you do this?

OpenStudy (anonymous):

(a-b)^2 - 14b^2 = \[((a-b)^2+ \sqrt{14}b)((a-b)^2+ \sqrt{14} b)\]

OpenStudy (anonymous):

sry did it incorrectly \[a ^{2}- 2ab + b ^{2} -16 b ^{2}\] = \[(a-b)^{2} -(4b)^{2}\] = \[(a - b -4b)(a-b+4b) = (a-5b)(a+3b)\]

OpenStudy (anonymous):

another method \[a ^{2}-5ab +3ab - 15b ^{2} = a(a-5b) -3b(a-5b)= (a-5b)(a-3b)\]

OpenStudy (anonymous):

does it help ???

OpenStudy (anonymous):

yes i see it now Thank youuu !!!!!

OpenStudy (anonymous):

u r welcome